OpenStudy (anonymous):

if y=14 when x=8 find y when x=12

Nnesha (nnesha):

is there any equation ??

OpenStudy (kittykitty04):

\[14=a*8\] \[\frac{ 14 }{ 2 }=\frac{ a*8 }{ 2 } \rightarrow 7=a*4 \rightarrow a= \frac{ 7 }{ 4 }\] so now we know \[a=\frac{ 7 }{ 4 }\] taking 14 divided by a will give you 8. now try 12 divided by a. you will get: \[12\div \frac{ 7 }{ 4 }=6\frac{ 6 }{ 7 }\]

OpenStudy (anonymous):

wait where did u get a from?